OpenStudy (anonymous):

verify (1/(sinx+1)) + (1/(cscx+1))=1

OpenStudy (anonymous):

(1 / sin(x) + 1) + (1/csc(x) + 1) put fractions over a common denominator: = [(csc(x) + 1) / (sin(x) + 1)(csc(x) + 1)] + [(sin(x) + 1) / (sin(x) + 1)(csc(x) + 1)] = [(csc(x) + 1) + (sin(x) + 1)] / [(sin(x) + 1)(csc(x) + 1)] = [csc(x) + sin(x) + 2] / [sin(x)csc(x) + sin(x) + csc(x) + 1] as csc(x) = 1/sin(x), we have: = [csc(x) + sin(x) + 2] / [sin(x)(1/sin(x)) + sin(x) + csc(x) + 1] = [csc(x) + sin(x) + 2] / [1 + sin(x) + csc(x) + 1] = [csc(x) + sin(x) + 2] / [sin(x) + csc(x) + 2] = 1 as required.

OpenStudy (anonymous):

thanks thats what i got